# A Survey on Deep Learning Hardware Accelerators for Heterogeneous HPC Platforms

### v2

@@ 4. Hardware Accelerators
Typical HPC workloads, like genomics, astrophysics, finance, and cyber security, require the elaboration of massive amounts of data and they can take advantage of DL methods with results that can surpass human ability (Bengio_2009, ; Schmidhuber_2015, ; Goodfellow_2016, ; SPAGNOLO20201, ). However, an ever-increasing computing power, a rapid change of the data analysis approaches, and the introduction of novel computational paradigms are needed. DL models rely on remarkable computational complexities that can be efficiently supported, without renouncing a good trade-off between speed, energy efficiency, design effort, and cost, by optimized hardware platforms that are able to provide high levels of parallelism and a considerable amount of memory resources. These platforms can be developed using CPUs, GPUs, FPGAs, CGRAs, and ASICs (Goodfellow_2016, ; DHILLESWARARAO_2022, ; Sze_2017, ; Machupalli_2022, ; Du_2018, ; Wang_2019, ; Leibo2019ACMSUR, ).
@@
Figure 2. Overview on state-of-the-art Neural Network accelerators based on available data collected in (Guo_Online, ).
Data collection available in (Guo_Online, ) provides a big picture on state-of-the-art hardware platforms suitable for Neural Network accelerators in terms of speed performances, energy consumption and energy efficiency. As it is well known, CPUs may have higher cache size and higher on-chip bandwidth than GPUs and reconfigurable architectures, but they show a limited ability to process large amounts of data in parallel. Consequently, as noticeable in Figure 2, CPU-based platforms are quite inefficient in comparison to their counterparts. On the other hand, with their high throughput and parallelism, GPUs are extremely efficient in terms of computational speed. Comparing the designs referenced in Figure 2 it can be appreciated that GPUs can process thousands of parallel threads and execute math and graphics tasks much more efficiently than general-purpose CPUs. Powerful GPUs are generally preferred for heavier tasks like training and running large, complex models built on large datasets. Conversely, FPGAs are well suited for AI applications that require customization to accelerate specific deep learning tasks that privilege energy efficiency over processing speeds. Speed performances achievable by heterogeneous computing platforms based on modern FPGAs are bounded to moderate levels by the limited amount of computing and memory resources available on chip (Qadeer2015, ; Sze_2017, ; DHILLESWARARAO_2022, ). Conversely, ASICs take longer design times, require higher design efforts, and do not offer flexibility, but they provide optimum computational speed and power consumption. A good trade-off between speed, power consumption, and design effort is offered by CGRAs that exhibit near-ASIC energy efficiency and performances with near-FPGA reconfigurability levels.
@@
Independently of the technology, a common problem in the design of accelerators is the energy and latency cost of accessing the off-chip DRAM memory, in particular considering the significant amount of data that the target HPC applications need to process. As sketched in Figure 3, different data reuse and stationary strategies can be exploited to reduce the number of accesses, each strategy offers a certain benefit (Park2015, ; Peemen2013, ; Sankaradas2009, ; Sriam2011, ; Cavigelli2017, ; Gupta2015, ). In weight stationary dataflow, convolutional weights are fixed and stored in the local memory of the Processing Elements (PEs) and reused on input activations uploaded step-by-step from the external DRAM. Conversely, in output stationary dataflow, partial outputs from the PEs are stored locally and reused step-by-step until the computation is completed. Then, just the final results are moved to the external DRAM. An efficient alternative is input stationary dataflow: in this case, input activations are stored in the PEs’ local memory, while weights are uploaded from the external DRAM and sent to the PEs. Another approach common to many accelerators is the introduction of quantization to reduce the data type width. Quantization represents an open challenge in designing DL models and many recent studies addressed this topic (quant2021, ; Liu2022, ). Integer or fixed-point data formats are generally preferred over the more computationally-intensive floating-point ones. This guarantees better memory occupation, lower computational cost, and higher robustness of the model (jin2022fnet, ). Extreme quantization techniques that use only one bit for the data stored (Binary Neural Networks (BNN2020, )) are widely used for very large networks but to get a comparable accuracy they require 2-11$\times$ the number of parameters and operations (Umuroglu2017, ), making them not suitable for complex problems.
@@ 4.1. Reconfigurable Hardware Accelerators
FPGAs and CGRAs are highly sought-after solutions to hardware accelerate a wide range of applications, including DL. The main feature of such reconfigurable platforms is the ability to support different computational requirements by repurposing the underlying hardware accelerators at deploy-time and also at runtime. More details on FPGA technologies and related EDA frameworks are respectively provided in Sections A.3 and A.4 of Appendix A. Several FPGA-based hardware accelerators for DL are structured as heterogeneous embedded systems (Ma2018, ; Yazdanbakhsh2018, ; Aimar2019, ; Perri2020, ; Li2021, ) that mainly consist of a general-purpose processor, for running the software workload; a computational module, designed to speed up common DL operators, like convolutions (Qiu2016, ; Venieris2019, ), de-convolutions (Chang2020, ; Sestito2021, ), pooling, fully connected operations, activation, and softmax functions (Spagnolo2022_1, ; Spagnolo2022_2, ); and a memory hierarchy to optimize data movement to/from the external DRAM to store data to be processed and computational results. A typical approach to accelerate convolutions consists of a systolic array architecture (SA), a regular pattern that can be easily replicated (Xuechao2017, ). Each PE in the array is a SIMD vector accumulation module where inputs and weights are supplied at each cycle by shifting them from the horizontally and vertically adjacent PEs (Figure 4(a)). The use of pipelined groups of PEs with short local communication and regular architecture enables a high clock frequency and limited global data transfer (Figure 4(b)).
Figure 4. FPGA accelerators: (a) Systolic array accelerator; (b) Pipelined dataflow accelerator.
@@
Although FPGAs have traditionally been proposed as accelerators for edge applications, they are starting to be adopted also in data centers. Microsoft’s Project Brainwave (Brainwave, ) uses several FPGA boards to accelerate the execution of RNNs in the cloud, exploiting the reconfigurability to adapt the platform to different DL models. One way to face the limitations imposed by the capability of FPGAs to effectively map very large DL models is to use a deeply pipelined multi-FPGA design. Recent studies focus on optimizing this type of architecture and maximizing the overall throughput (Zhang2016, ; Rahman2017, ; Junnan2019, ). In these application contexts, CGRAs represent an alternative to FPGAs, providing reconfigurability with coarser-grained functional units. They are based on an array of PEs, performing the basic arithmetic, logic, and memory operations at the word level and using a small register file as temporary data storage. Neighboring PEs are connected through reconfigurable routing that allows transferring intermediate results of the computations towards the proper neighbors for the next computational step. CGRAs can represent a powerful solution to accelerate dense linear algebra applications, such as ML, image processing, and computer vision (amber2022vlsi, ; tangram, ). Thanks to parallel computing and time-multiplexing, CGRAs can efficiently support and combine spatial and temporal computational models. Furthermore, they are flexible enough for specific domains, and their interconnections, being not as complex as those present on FPGAs, provide remarkable advantages in terms of speed, energy efficiency, and resource utilization.
@@ 4.2. ASIC-based Accelerators
To comply with the computational capabilities required by AI workloads, new powerful processing architectures are upcoming. Among them, there are two different types of Neural Processing Units: single-chip NPUs and NPUs integrated in the general purpose CPU. One of the main trends toward the next generation of laptops follows the second option by pushing the performance of AI workloads by integrating into the general-purpose CPU not only a GPU to accelerate graphics but also an NPU. This is the case the recent Lunar Lake Intel processor architecture (lunarlake, ). Table 4 is an attempt to offer a common ground of different types of AI-accelerators in terms in terms of process technology node, area, peak performance, energy efficiency and maturity level.
The purpose of an integrated Neural Processing Unit is to accelerate the performance and improve the energy efficiency of specific AI-tasks offloaded from the CPU (song2019isscc, ). In particular, NPUs are designed to accommodate a reasonable amount of multiply/accumulate (MAC) units, which are the PEs devised in the convolutional and fully-connected layers of DNNs (Chen_2017, ; desoli17isscc, ).
@@ 4.3. Accelerators for Sparse Matrices
This section overviews accelerator architectures designed to take advantage of computations suitable to sparse matrices. Definitions, storage formats appropriate for sparse matrices, and their impacts on the computational complexity of DL models are discussed in Appendix A.6.
Some accelerators (e.g., Cnvlutin, Cambricon-X, Eyeriss) handle only one-sided sparsity, which stems either from zero-valued activations or network pruning, thus achieving only a partial reduction in compute and data reduction. On the other hand, other accelerators (e.g., SCNN, SparTen, Eyeriss v2) target two-sided sparsity, which originates from network pruning and zero-valued activations. In addition to the different approaches to exploiting sparsity, these architectures also employ distinct dataflows to execute the DNN layers. Due to the complexity of the logic, existing hardware accelerators for sparse processing are typically limited to a specific layer type (e.g., fully-connected layers, convolutional layers).
@@
Eyeriss (Chen_2017, ) targets CNN acceleration by storing in DRAM only nonzero-valued activations in Compressed Sparse Columns (CSC) format and by skipping zero-valued activations (by means of gating the datapath switching and memory accesses) to save energy. Eyeriss v2 (eyerissv2:2019, ), which targets DNNs on mobile devices, also supports sparse network models. It utilizes the CSC format to store weights and activations, which are kept compressed not only in memory but also during processing. To improve flexibility, it uses a hierarchical mesh for the PEs interconnections. By means of these optimizations, Eyeriss v2 is significantly faster and more energy-efficient than the original Eyeriss.
@@
Cnvlutin (cnvlutin:isca2016, ), which also targets CNN acceleration, uses hierarchical data-parallel units, skips computation cycles for zero-valued activations and employs a co-designed data storage format based on Compressed Sparse Rows (CSR) to compress the activations in DRAM. However, it does not consider the sparsity of the weights. On the contrary, Cambricon-X architecture (zhang2016micro, ) exploits the sparsity of CNNs by enabling the PEs to store the compressed weights in CSR format for asynchronous computation. However, it does not exploit activation sparsity. EIE (han2016isca, ), besides compressing the weights through a variant of CSC sparse matrix representation and skipping zero-valued activations, employs a scalable array of PEs, each storing a partition of the DNN in SRAM that allows obtaining significant energy savings with respect to DRAM. NullHop (Aimar2019, ) is a CNN accelerator architecture that applies the Compressed Image Size (CIS) format to the weights and skips the null activations, similarly to EIE. Sparse CNN (SCNN) (SCNN:isca2017, ) is an accelerator architecture for inference in CNNs. It employs a cluster of asynchronous PEs comprising several multipliers and accumulators. SCNN exploits sparsity in both weights and activations, which are stored in the classic CSR representation. It employs a Cartesian product-based computation architecture that maximizes the reuse of weights and activations within the cluster of PEs; the values are delivered to an array of multipliers, and the resulting scattered products are summed using a dedicated interconnection mesh. By exploiting two-sided sparsity, SCNN improves performance and energy over dense architectures. SparTen (sparten:micro2019, ) is based on SCNN (SCNN:isca2017, ). It addresses some considerable overheads of SCNN in performing the sparse vector-vector dot product by improving the distribution of the operations to the multipliers and allows using any convolutional stride (not being limited to unit-stride convolutions as SCNN). It also addresses unbalanced sparsity distribution across the PEs employing an offline software scheme. The PermDNN architecture (perm:micro2018, ) addresses the generation and execution of hardware-friendly structured sparse DNN models using permuted diagonal matrices. In this way, it does not incur load imbalance which is caused by the irregularity of unstructured sparse DNN models.
@@
SqueezeFlow (squeezeflow:2019, ) is an accelerator architecture that exploits the sparsity of CNN models. To reduce hardware complexity, it exploits concise convolution rules to benefit from the reduction of computation and memory accesses as well as the acceleration of existing dense CNN architectures without intrusive PE modifications. The Run Length Compression (RLC) format is used to compress activations and weights. A different strategy is pursued by the Unique Weight CNN (UCNN) accelerator (ucnn:isca2018, ), which proposes a generalization of the sparsity problem. Rather than considering only the repetition of zero-valued weights, UCNN exploits repeated weights with any value by reusing CNN sub-computations and reducing the model size in memory. SIGMA (sigma:2020, ) is an accelerator for DNN training, which is characterized by a flexible and scalable architecture that offers high utilization of its PEs regardless of kernel shape (i.e., matrices of arbitrary dimensions) and sparsity pattern. It targets the acceleration of GEMMs with unstructured sparsity. Bit-Tactical (bittactical:2019, ) is a DNN accelerator where the responsibility for exploiting weight sparsity is shared between a static scheduling middleware and a co-designed hardware front-end, with a lightweight sparse shuffling network which comprises two multiplexers per activation input. Unlike SIGMA and other accelerators, Bit-tactical leverages scheduling in software to align inputs and weights.
Flexagon (flexagon:aplos2023, ) is a reconfigurable accelerator capable of performing sparse-sparse matrix multiplication computation by using the particular data flow that best matches each case.
@@
Besides the design of specialized hardware accelerators to exploit model sparsity, a parallel trend is to use GPU architectures. Pruned sparse models with unstructured sparse patterns introduce irregular memory accesses that are unfriendly on commodity GPU architectures. The first direction to tackle this issue on commodity DNN accelerators is at the software layer, using pruning algorithms that enforce a particular sparsity pattern, such as tile sparsity (guo:sc2020, ), on the model that allows leveraging existing GEMM accelerators. A second direction to leverage the sparsity in DNN models on GPUs is to introduce new architectural support, such as Sparse Tensor Cores (sparsetensorcore:2021, ). The NVIDIA Ampere architecture introduces this Sparse Tensor Core design with a fixed 50% weight pruning target and achieves a better accuracy and performance trade-off. However, sparsity from activations, which are dynamic and unpredictable, is challenging to leverage on GPUs. Indeed, the current Sparse Tensor Core is only able to take advantage of weight sparsity but not activation sparsity. Reconfigurability appears to be a keyword for the design of new sparse accelerators because some network models exhibit dynamic sparsity (fedus:jmlr2022, ), where the position of non-zero elements changes over time.

### v3

@@ 4. Hardware Accelerators
Typical HPC workloads, like genomics, astrophysics, finance, and cyber security, require the elaboration of massive amounts of data and they can take advantage of DL methods with results that can surpass human ability (Bengio_2009, ; Schmidhuber_2015, ; Goodfellow_2016, ; SPAGNOLO20201, ). However, an ever-increasing computing power, a rapid change of the data analysis approaches, and the introduction of novel computational paradigms are needed. DL models can be efficiently supported by optimized hardware platforms providing high levels of parallelism and a considerable amount of memory resources. These platforms can be developed using CPUs, GPUs, FPGAs, and ASICs (Goodfellow_2016, ; DHILLESWARARAO_2022, ; Sze_2017, ; Machupalli_2022, ; Du_2018, ; Wang_2019, ; Leibo2019ACMSUR, ).
@@
Figure 2. Overview on state-of-the-art Neural Network accelerators based on available data collected in (Guo_Online, ). Legenda: Simulation means GOPS/W values collected from post-layout simulation; Test means from prototype devices; Product means from off-the-shelf devices.
Figure 2 presents a comparison of state-of-the-art architectures in terms of speed (Giga Operations per Second) versus power consumption (Watt). The diagonal dashed lines represent energy efficiency levels in (GOPS/W) and TOPS/W (Tera Operations per Second per Watt), with higher slopes indicating better energy efficiency. The most energy-efficient devices are ASICs and GPUs clustering in the high range of energy efficiency (1–100 TOPS/W) and mainly located in the top right region characterized by the highest computational throughput. Powerful GPUs are generally preferred for heavier tasks like training and running large, complex models built on large datasets. Conversely, FPGAs are well suited to accelerate specific inference tasks that privilege lower power consumption over processing speeds. From the figure, FPGAs (represented in blue) are well distributed across different power and performance levels, mostly clustering in the lower-to-mid range of energy efficiency (from 10 GOPS/W to 1 TOPS/W).
Independently of the technology, a common problem in the design of accelerators is the energy and latency cost of accessing the off-chip DRAM memory, in particular considering the significant amount of data that HPC applications need to process. As sketched in Figure 3, different data reuse and stationary strategies can be exploited to reduce the number of accesses (Park2015, ; Peemen2013, ; Sankaradas2009, ; Sriam2011, ; Cavigelli2017, ; Gupta2015, ). In weight stationary dataflow, convolutional weights are fixed and stored in the local memory of the Processing Elements (PEs) and reused on input activations uploaded step-by-step from the external DRAM. Conversely, in output stationary dataflow, partial outputs from PEs are stored locally and reused step-by-step until the computation is completed. Then, just the final results are moved to the external DRAM. An efficient alternative is input stationary dataflow: input activations are stored in the PEs’ local memory, while weights are uploaded from the external DRAM and sent to the PEs. Another approach common to many accelerators is the introduction of quantization to reduce the data type width (quant2021, ; Liu2022, ). Integer or fixed-point data formats are generally preferred to the more computationally intensive floating-point ones. This guarantees better memory occupation and lower computational cost (jin2022fnet, ). Extreme quantization techniques that use only one bit for the data stored (Binary Neural Networks (BNN2020, )) are widely used for very large networks but to get a comparable accuracy they require 2-11$\times$ the number of parameters and operations (Umuroglu2017, ), making them not suitable for complex problems.
@@ 4.1. Reconfigurable Hardware Accelerators
FPGAs and Coarse-Grained Reconfigurable Arrays (CGRAs) are highly sought-after solutions to hardware accelerate a wide range of applications. The main feature of such platforms is the ability to support different computational requirements by repurposing the underlying hardware accelerators at deploy-time and also at runtime. More details on FPGA technologies and related EDA frameworks are respectively provided in Sections A.3 and A.4 of Appendix A. Several FPGA-based hardware accelerators for DL are structured as heterogeneous embedded systems (Ma2018, ; Yazdanbakhsh2018, ; Aimar2019, ; Perri2020, ; Li2021, ) that mainly consist of a general-purpose processor, for running the software workload; a computational module, designed to speed up common DL operators, like convolutions (Qiu2016, ; Venieris2019, ), de-convolutions (Chang2020, ; Sestito2021, ), pooling, fully connected operations, activation, and softmax functions (Spagnolo2022_1, ; Spagnolo2022_2, ); and a memory hierarchy to optimize data movement to/from the external DRAM to store data to be processed and computational results. A typical approach to accelerate convolutions consists of a systolic array architecture (SA), a regular pattern that can be easily replicated (Xuechao2017, ). Each PE in the array is a SIMD vector accumulation module where inputs and weights are supplied at each cycle by shifting them from the horizontally and vertically adjacent PEs. The use of pipelined groups of PEs with short local communication and regular architecture enables a high clock frequency and limited global data transfer.
@@
Although FPGAs have traditionally been proposed as accelerators for edge applications, they are starting to be adopted also in data centers. Microsoft’s Project Brainwave (Brainwave, ) uses several FPGA boards to accelerate the execution of RNNs in the cloud, exploiting the reconfigurability to adapt the platform to different DL models. One way to face the limitations imposed by the capability of FPGAs to effectively map very large DL models is to use a deeply pipelined multi-FPGA design. Recent studies focus on optimizing this type of architecture and maximizing the overall throughput (Zhang2016, ; Rahman2017, ; Junnan2019, ). In these application contexts, CGRAs represent an alternative to FPGAs, providing reconfigurability with coarser-grained functional units. They are based on an array of PEs, performing the basic arithmetic, logic, and memory operations at the word level and using a small register file as temporary data storage. Neighboring PEs are connected through reconfigurable routing that allows transferring intermediate results of the computations towards the proper neighbors for the next computational step. CGRAs can represent a vaible solution to accelerate dense linear algebra applications, such as ML, image processing, and computer vision (amber2022vlsi, ; tangram, ).
@@ 4.2. ASIC-based Accelerators
To comply with the computational capabilities required by AI workloads, new powerful processing architectures are upcoming. Among them, there are two different types of Neural Processing Units: single-chip NPUs and NPUs integrated in the general purpose CPU. One of the main trends toward the next generation of laptops follows the second option by pushing the performance of AI workloads by integrating into the general-purpose CPU not only a GPU to accelerate graphics but also an NPU. This is the case of the recent Lunar Lake Intel processor architecture (lunarlake, ). Table 4 is an attempt to offer a common ground of different types of AI-accelerators in terms of process technology node, area, peak performance, energy efficiency and maturity level.
The purpose of an integrated NPU is to accelerate the performance and improve the energy efficiency of specific AI-tasks offloaded from the CPU (song2019isscc, ). In particular, NPUs are designed to accommodate a reasonable amount of multiply/accumulate (MAC) units, which are the PEs devised in the convolutional and fully-connected layers of DNNs (Chen_2017, ; desoli17isscc, ).
@@ 4.3. Accelerators for Sparse Matrices
Network pruning and zero-valued activations introduce sparsity that can be exploited by hardware accelerators to achieve compute and data reduction. This section overviews accelerator architectures designed to manage sparse matrices. Definitions, storage formats appropriate for sparse matrices, and their impacts on the computational complexity of DL models are discussed in Appendix A.6.
@@
Eyeriss (Chen_2017, ) targets CNN acceleration by storing in DRAM only nonzero-valued activations in Compressed Sparse Columns (CSC) format and by skipping zero-valued activations to save energy. Eyeriss v2 (eyerissv2:2019, ), which targets DNNs on mobile devices, also supports sparse network models. It utilizes the CSC format to store weights and activations, which are kept compressed not only in memory but also during processing. To improve flexibility, it uses a hierarchical mesh for the PEs interconnections. By means of these optimizations, Eyeriss v2 is significantly faster and more energy-efficient than the original Eyeriss.
@@
Cnvlutin (cnvlutin:isca2016, ) uses hierarchical data-parallel units, skips computation cycles for zero-valued activations and employs a co-designed data storage format based on Compressed Sparse Rows (CSR) to compress the activations in DRAM. However, it does not consider the sparsity of the weights. On the contrary, Cambricon-X architecture (zhang2016micro, ) enables the PEs to store the compressed weights in CSR format for asynchronous computation. However, it does not exploit activation sparsity. EIE (han2016isca, ), besides compressing the weights through a variant of CSC sparse matrix representation and skipping zero-valued activations, employs a scalable array of PEs, each storing a partition of the DNN in SRAM that allows obtaining significant energy savings with respect to DRAM. NullHop (Aimar2019, ) applies the Compressed Image Size (CIS) format to the weights and skips the null activations, similarly to EIE. Sparse CNN (SCNN) (SCNN:isca2017, ) is an accelerator architecture for inference in CNNs. It employs a cluster of asynchronous PEs comprising several multipliers and accumulators. SCNN exploits sparsity in both weights and activations, which are stored in the classic CSR representation. It employs a Cartesian product-based computation architecture that maximizes the reuse of weights and activations within the cluster of PEs; the values are delivered to an array of multipliers, and the resulting scattered products are summed using a dedicated interconnection mesh. By exploiting two-sided sparsity, SCNN improves performance and energy over dense architectures. SparTen (sparten:micro2019, ) is based on SCNN (SCNN:isca2017, ). It addresses some considerable overheads of SCNN in performing the sparse vector-vector dot product by improving the distribution of the operations to the multipliers and allows using any convolutional stride. It also addresses unbalanced sparsity distribution across the PEs employing an offline software scheme. The PermDNN architecture (perm:micro2018, ) uses permuted diagonal matrices to not generate load imbalance which is caused by the irregularity of unstructured sparse DNN models.
@@
SqueezeFlow (squeezeflow:2019, ) exploits concise convolution rules to benefit from the reduction of computation and memory accesses as well as the acceleration of existing dense CNN architectures without intrusive PE modifications. The Run Length Compression (RLC) format is used to compress activations and weights. A different strategy is pursued by the Unique Weight CNN (UCNN) accelerator (ucnn:isca2018, ), which proposes a generalization of the sparsity problem. Rather than considering only the repetition of zero-valued weights, UCNN exploits repeated weights with any value by reusing CNN sub-computations and reducing the model size in memory. SIGMA (sigma:2020, ) is characterized by a flexible and scalable architecture that offers high utilization of its PEs regardless of kernel shape (i.e., matrices of arbitrary dimensions) and sparsity pattern. It targets the acceleration of GEMMs with unstructured sparsity. Bit-Tactical (bittactical:2019, ) uses a static scheduling middleware and a co-designed hardware front-end, with a lightweight sparse shuffling network that comprises two multiplexers per activation input. Unlike SIGMA and other accelerators, Bit-tactical leverages scheduling in software to align inputs and weights. Flexagon (flexagon:aplos2023, ) is a reconfigurable accelerator capable of performing sparse-sparse matrix multiplication computation by using the particular data flow that best matches each case.
@@
Besides the design of specialized hardware accelerators to exploit model sparsity, a parallel trend is to use GPU architectures. Pruned sparse models with unstructured sparse patterns introduce irregular memory accesses that are unfriendly on commodity GPU architectures. The first direction to tackle this issue is at the software layer, using pruning algorithms that enforce a particular sparsity pattern, such as tile sparsity (guo:sc2020, ), on the model that allows leveraging existing GEMM accelerators. A second direction is to introduce new architectural support, such as Sparse Tensor Cores (sparsetensorcore:2021, ). The NVIDIA Ampere architecture introduces this design with a fixed 50% weight pruning target and achieves a better accuracy and performance trade-off. However, sparsity from activations, which are dynamic and unpredictable, is challenging to leverage on GPUs. Indeed, the current Sparse Tensor Core can take advantage of weight sparsity. Reconfigurability appears to be a keyword for the design of new sparse accelerators because some network models exhibit dynamic sparsity (fedus:jmlr2022, ), where the position of non-zero elements changes over time.